# SocialEyes: Scaling mobile eye-tracking to multi-person social settings

### v3

@@ 4.1.3. Stimuli
Music Performance: The hour-long composition consisted of home-made instruments, spoken texts, and electronic soundscapes to present the audience with dramatic soundscapes that explore issues related to the American criminal justice system. Of the 17 total tableaux included in the performance, some were too long, uncomfortable, confusing; others were direct, melodic, and in familiar styles.
Film:The film showcases the compositional process behind the performance, engaging with exonerees, whose story and reaction to the performance is also captured.
@@ 4.1.4. Study Design
The current utility test takes place within the context of a larger project examining the role of context information in altering audience physiological responses and social attitudes and behaviours.During the event, in addition to wearing eye-tracking glasses, participants also wore a smart watch to monitor cardiac activity and completed a series of surveys. Parallel to this in-person event, we also presented a livestream online. We recorded eye measures, via web browser and webcam ((Saxena et al., 2022, 2023)). We were interested in the effect social co-presence might have on our measures of interest. Detailed analyses related to our cognitive neuroscientific and social psychological questions will be reported elsewhere. In the current study, we focus on validating the utility of our software framework–SocialEyes.
@@ 4.1.5. Procedure
Before the event, ADB connection was established between the server and each smartphone device. A forced synchronisation to the assigned NTP server on the smartphones was also done. Participants were seated evenly across rows two to four in the audience, hereby referred to as participant-rows one to three in this paper. After taking their seat, participants were instructed on how to put the glasses on. Trained research assistants completed a quick calibration procedure. Note that explicit eye-egoview calibration was not needed for the eye-tracking hardware (see 4.1.1). The calibration procedure mentioned here simply performs an individual offset correction using a single fixation target, which can account for physiological factors, such as the Kappa angle between the optical and visual axis of the eye. Participants retained their seating positions over the entire study duration.
To ensure proper connection to all eye-tracking devices, test recordings for a short length (1-5 minutes) were made before the actual recording session was started. We started the recording session before the event host came on stage. After the start of the event, no interruptions were allowed and all interactions with the recording devices were done remotely using the GlassesRecord module TUI to trigger recordings and monitor the eye-tracking devices (see GlassesRecord in 4.1.2).
@@ 4.2.1. Time Synchronisation
Offsets results
Offsets: Time offsets and mean roundtrip duration between each device and the server was recorded at regular time intervals of 10 seconds during each of the four sessions. The mean offset over all devices was 19.58 ms (SD = 8.50 ms) for G1F, 45.57 ms (SD = 16.45 ms) for G1P, 18.52 ms (SD = 8.46 ms) for G2P, and 37.65 ms (SD = 16.14 ms) for G2F; see Figure 4A. Mean offsets were computed after outlier rejection of 3 data points that had high offsets (-277.4 ms, -261.4 ms, and 277.3 ms) due to a failure in the initial time synchronisation procedure for the respective devices. The higher offset of the outliers, however, does not hinder their time synchronisation with other devices since the relative offset as well as the over-time drift of the offset were recorded. The mean roundtrip duration over all 30 devices, was 0.51 ms (SD = 0.04 ms) for G1F, 0.52 ms G1P (SD = 0.03 ms), 0.49 ms (SD = 0.03 ms) for G2P, and 0.53 ms (SD = 0.02 ms) for G2F; see Figure 5B.
Drifts: Accumulated drifts in time offsets were calculated as the difference between the offset at a given time point from the initial offset. The mean drift over all 30 devices for G1F was 8.58 ms (SD = 6.48 ms); G1P: 8.48 ms (SD = 5.54 ms); G2P: 11.52 ms (SD = 6.53 ms); G2F: 9.92 ms (SD = 6.88 ms).
RANSAC Fit: The mean accuracy score of the linear fits of the 30 devices for each session was 0.95 (SD = 0.03) for G1F, 0.95 (SD = 0.04) for G1P, 0.95 (SD = 0.06) for G2P, and 0.95 (SD = 0.04) for G2F.
@@ 4.2.2. Visualisation
Visualisations
Raw egocentric gaze coordinates and homography-transformed gaze coordinates were projected onto each viewer’s egoview and the shared centralview, respectively, after removal of blinks. The visualisations (presented in Figure 5) enable data validation and demonstrate accurate projective transformation using the SuperGlue features identified in the Homography module (see 3.2.6). The transforms computed using our SuperGlue + RANSAC approach handle video artefacts, such as motion blur (from sudden head movements), poor lighting/texture (frames with large dark regions, shadows, lens flare, glare etc.), obstructions (audience members in previous rows), and scale variations (different resolution and zoom in centralview and egoviews) reliably. However, further optimisations and investigations would be needed to identify edge cases in the homography estimation (we discuss these optimisations in 5.1.1). Figure 5 displays still, single frame examples of our transformed data; an example video can be viewed here: https://tinyurl.com/multipersonET. In the first part of this video example (tableau: “canjo”), the two performers are seated very near to each other. Even at this close distance, we can clearly distinguish audience-gaze on one or the other performer, vs. their instruments, and follow shifts in audience attention between these elements. When this “canjo” tableau ends, the two performers spread out and we see the gaze follow them around the stage. Once one of the performers starts speaking (tableau: “pod rattle incident”); the audience mostly maintain their focus on him, rarely glancing toward the other performer who is gently striking metallic objects. This video highlights the observation that the transformed gaze reliably follows the two musicians around the stage and provides proof-of-concept that our homography approach successfully projects gaze from a large audience in challenging scenes.
@@
To study the gaze patterns over an evolving scene and compare gaze data between different viewers, we map gaze from viewer’s frame of reference to the scene’s frame of reference with our Homography module (see 3.2.6). Figure 7 presents individual gaze heatmaps from all participants and the aggregate gaze activity from all participants for each session overlaid on a reference frame from that session. Projecting the gaze to the static centerview perspective gets rid of the independent positions and head movements of each participant, thereby getting rid of the central bias. Heatmaps of transformed gaze coordinates (Figure 7) demonstrate higher gaze dispersion and multiple discrete clusters for each session duration. These clusters coincide with the presentation screen area for the film viewing sessions and the performers’ movement activity for the concert performance sessions (see averaged heatmaps for each session in Figure 7). Higher gaze dispersion can be observed during performance viewing than during film viewing, on both days, which could be partially explained by the larger physical space covered on the stage by the performers during the concert, in comparison to the presentation screen-size for the film.
Egocentric heatmaps
Transformed heatmaps
Averaged heatmaps
Investigating the gaze heatmaps for different participant seating clusters, we observe that participants in the right half of the audience spent more time exploring the right side of the stage, reflected by larger (higher gaze dispersion) and more intense (higher gaze duration) clusters in the right half and vice versa (see Figure 8A; heatmaps are arranged by seating order). This effect is, however, only observed in the performance sessions and not when the participants viewed the pre-recorded documentary film. The heatmaps could further be aggregated to observe similar effects with different participant clusters, such as clustering by participant columns (Figure 8A) or rows (Figure 8B). Figure 8 highlights different trends in the position of the heatmap centroid (i.e., the peak heatmap intensity location; marked by the intersection of white dashed lines), based on seating position.
@@ 4.2.3. Analysis
Further comparisons between individual heatmaps can be made with the help of metrics from our Analysis module (see 3.2.8). Pairwise SIM and CC scores were calculated for each participant pair for each of the four event sessions. For raw egocentric gaze, the mean SIM score during film viewing was 0.35 (SD = 0.21) for G1F and 0.38 (SD = 0.19) for G2F, while the mean SIM during the performance was 0.40 (SD = 0.18) for G1P and 0.43 (SD = 0.17) for G2P. The mean CC was 0.37 (SD = 0.30) for G1F, 0.42 (SD = 0.28) for G2P, 0.43 (SD = 0.26) for G1P and 0.46 (SD = 0.25) for G2P. When using homography-transformed gaze, the scores in all sessions, for both of these metrics, increase drastically (see Figure 9; left column), nearly doubling in most cases. A 2D chart of seat positions (row, column) for each participant (see Figure 6 and 7 for audience seating positions on each day) was used to calculate euclidean distances (DI) between each participant pair. Calculated SIM scores on the homography-transformed gaze were found to be negatively correlated with the euclidean distances during the performance session on both days (i.e., the closer two participants were to each other, the stronger was the correlation between their gaze patterns and vice versa). The Pearson correlation coefficient between SIM and DI was -0.10 (p = 0.039) for G1P and -0.17 (p ¡ 0.001) for G2P.
@@
The mean gaze entropy over all participants during film viewing was 0.76 (SD = 0.03) for G1F and 0.77 (SD = 0.03) for G2F. The mean entropy was higher when participants watched the performance, with a value of 0.82 (SD = 0.03) for G1P and 0.83 (SD = 0.03) for G2P. The higher values represent more spatial gaze dispersion when participants watch the performance as compared to the documentary film. The average gaze velocity (calculated in pixels/sample) during film viewing was found to be 22.99 (SD = 7.25) for G1P and 36.00 (SD = 5.90) for G2P. During the performance sessions, the average gaze velocity was 16.76 (SD = 6.79) for G1P and 22.11 (SD = 8.03) for G2P. Interestingly, gaze velocity was higher during film viewing on both days which signifies longer saccadic movements when viewing pre-recorded media on a screen, in comparison to a live concert performance
Spatial gaze measures
Blinks
@@ 4.2.4. Exploratory Temporal Analysis
The synchronised gaze data from multiple devices mapped to a common coordinate space (centralview) further facilitates novel spatiotemporal investigations of collective gaze measures over time. In Figure 11, we visualise the measures of standard deviation, minimum convex area, and points in frame, over the entire duration of a film-viewing session (1 hr 20 min film, plus some minutes before/after) from 30 people (G1). These measures–computed by the Analysis module (see 3.2.8)–reveal collective gaze dynamics over time. Standard deviation of the gaze distribution of all participants at any given time is represented in the horizontal and vertical directions by the ”SD x” and ”SD y” measures (top and second panels, respectively). The ”Normalized convex area” (third panel) represents overall gaze dispersion as the area of a minimum convex hull enclosing gaze from all participants. The number of gaze points falling within the centralview scene at a given timepoint is represented by the ”Points in Frame” measure (bottom panel). The time series of all four measures highlight clear event boundaries such as the start and stop of the film presentation (vertical dotted lines). Future analyses will explore which features of the film drive changes in these collective gaze dynamics.
Temporal metrics
In another example, visualising the Normalised Convex Area measure for the entire concert performance (Figure 12) shows similar collective gaze behaviour from audiences during each performance tableau (represented by vertical colored lines) for each participant group. These collective gaze measures and corresponding time series could be further analysed to investigate temporal interactions with the audiovisual stimulus features (outside the scope of the current analysis). We provide these examples here to highlight the rich and vast amount of possibilities from shared gaze metrics in dynamic social settings.
Contour area analysis
@@ 5. General Discussion
Results from our utility test demonstrate the end-to-end implementation of the recording mode of SocialEyes in a real-world application. SocialEyes tackles 1) critical challenges of time-synchronisation of multi-sensor data to scale data collection, 2) monitoring and troubleshooting to ensure reliable data collection over long durations, 3) semantic gaze mapping of the synchronised gaze data to analyse collective gaze patterns, and finally, 4) shared gaze visualisation and analysis to achieve novel insights from the collected data. The proposed framework aims to reduce technological and logistical barriers in conducting naturalistic eye-tracking experiments with large social groups, allowing replication/refutation and extension of previous lab studies in social, naturalistic contexts (Kulke and Hinrichs, 2021; Laidlaw et al., 2016; Foulsham and Kingstone, 2017).
@@ 5.1.1. Edge-cases in gaze projection
Future objectives for our framework would be to further investigate and optimise gaze mapping accuracy in different environments. In our previous pilot study (Saxena and Fink, 2023), we identified a few edge case scenarios that pose a challenge to the homography transformations. In summary, the SuperGlue + RANSAC approach demonstrated robustness to head movements, illumination changes, and scene occlusions, as observed in our current utility study as well. However, close distances between the observer and scene as well as duplicate objects in a scene can result in false predictions. Systematically identifying these edge failure cases in challenging environments and addressing them—either through software and algorithmic updates or by establishing well-defined best practices for experimental setups—could significantly improve the robustness and reliability of gaze projections. To automate the collection of ground-truth gaze locations in these investigations, fiducial markers can be used by guiding participants to fixate on specific targets––as demonstrated in Appendix A.
@@ 5.1.2. Alternatives to planar homography
The assumption of a substantial distance between the observer(s) and the scene is a theoretical limitation of the current homography-based gaze projection as it projects a plane at sufficiently large (infinite) distance from one view to another. This assumption allows for more reliable and calibration-free projections, particularly when dealing with moving cameras, such as the head-mounted egoview, and dynamic (non-static) scenes, such as a music performance. We show the validity of this assumption for our current setup in appendix A. Future optimisations can also consider multi-view homography where the scene is captured by multiple centralview cameras and each egoview camera is related with the centralview corresponding to its position. This strategy would enable more precise cross-view analyses and spatial consistency in wide-view environments. To eliminate the 2D plane assumption, two alternative approaches are possible. Given the intrinsic parameters of the two cameras (egoview and centralview) and a common 3D coordinate system, a 2D point from the egoview can be projected to the 3D world coordinate system and then to the 2D centralview. This approach relies on pre-existing information about the locations of objects in the 3D world space, which is highly unfeasible, particularly in dynamic scenes. The second approach involves calibrating extrinsic positions of each egoview-centralview camera pair and projecting the 2D point from one camera’s view to another using epipolar geometry. This triangulation approach provides a general (ideal) solution, particularly for scenes with varying depths or close viewing distances where the homography planar assumption is not valid. However, it relies on accurate calibration parameters between the two cameras which is challenging with non-stationary cameras such as the head-mounted egoview cameras. As a result, the application of this approach relies on accurate and frequent re-calibration of the egoview cameras and prominent (unobstructed, well-lighted, strategically placed) visual markers that can be detected in the dynamic scene. These requirements might often interfere with the experimental setup, since the control of lighting and marker placement is not always possible. Additionally, markers create a visual distraction for viewers that would interfere with scientific investigations of visual attention. As an alternative to visual markers, calibration with invisible infrared markers (Dogan et al., 2022) or pose estimation of the moving egoview camera using algorithms such as visualSLAM (Chen et al., 2022) and optical flow (Beauchemin and Barron, 1995) is also possible. A potentially more direct approach would be to infer the pose of the egoview camera at each timepoint by using Inertial Measurement Unit (IMU) sensors. The eye-tracking hardware used in our utility study (see 4.1.1) are equipped with on-device 9DOF IMU. However, as mentioned previously, the IMU stream had to be disabled in favour of gaze data reliability. Overall, triangulation based methods might offer benefits, such as depth compensation and runtime efficiency, warranting future investigations to evaluate their tradeoffs in performance, accuracy, and ecological validity compared to the current homography approach.

### v4

@@ 4.1.3. Stimuli
[…]
@@ 4.1.4. Study Design
The current utility test takes place within the context of a larger project examining the role of context information in altering audience physiological responses and social attitudes and behaviours. During the event, in addition to wearing eye-tracking glasses, participants also wore a smart watch to monitor cardiac activity and completed a series of surveys. Parallel to this in-person event, we also presented a livestream online. We recorded eye measures, via web browser and webcam (Saxena et al., 2022, 2024). We were interested in the effect social co-presence might have on our measures of interest. Detailed analyses related to our cognitive neuroscientific and social psychological questions will be reported elsewhere. In the current study, we focus on validating the utility of our software framework–SocialEyes.
@@ 4.1.5. Procedure
Before the event, ADB connection was established between the server and each smartphone device. A forced synchronisation to the assigned NTP server on the smartphones was also done. Participants were seated evenly across rows two to four in the audience, hereby referred to as participant-rows one to three in this paper. After taking their seat, participants were instructed on how to put the glasses on. Trained research assistants completed a quick calibration procedure. Note that explicit eye-egoview calibration was not needed for the eye-tracking hardware (see Section 4.1.1). The calibration procedure mentioned here simply performs an individual offset correction using a single fixation target, which can account for physiological factors, such as the Kappa angle between the optical and visual axis of the eye. Participants retained their seating positions over the entire study duration.
To ensure proper connection to all eye-tracking devices, test recordings for a short length (1-5 minutes) were made before the actual recording session was started. We started the recording session before the event host came on stage. After the start of the event, no interruptions were allowed and all interactions with the recording devices were done remotely using the GlassesRecord module TUI to trigger recordings and monitor the eye-tracking devices (see GlassesRecord in Section 4.1.2).
@@ 4.2.1. Time Synchronisation
[Time synchronisation results]Figure 4. A. Mean time offsets are presented on a y-axis range of 0-80 ms. Mean offsets in the second recording session is higher than the first session on both days. Mean Roundtrip durations are presented in the right plot on a y-axis range of 0 - 0.575 ms. B. Drifts in offset for each device is plotted on a y-axis range of 0-30 ms. Spread of the recorded drifts vary slightly between devices but the recorded drifts do not contain spurious outliers.
[…]
@@ 4.2.2. Visualisation
[Visualisations from SocialEyes]Figure 5. Fully described in the text.
[Egocentric gaze heatmaps]Figure 6. Fully described in the text.
[Transformed gaze heatmaps]Figure 7. Fully described in the text.
[Averaged heatmaps]Figure 8. Fully described in the text.
Egocentric and homography-transformed gaze coordinates were projected onto each viewer’s egoview and the shared centralview, respectively, after removal of blinks. The visualisations (presented in Figure 5) enable data validation and demonstrate accurate projective transformation using the SuperGlue features identified in the Homography module (see Section 3.2.6). The transforms computed using our SuperGlue + RANSAC approach handle video artefacts, such as motion blur (from sudden head movements), poor lighting/texture (frames with large dark regions, shadows, lens flare, glare etc.), obstructions (audience members in previous rows), and scale variations (different resolution and zoom in centralview and egoviews) reliably. However, further optimisations and investigations would be needed to identify edge cases in the homography estimation (we discuss these optimisations in Section 5.1.1). Figure 5 displays still, single frame examples of our transformed data; an example video can be viewed here: https://tinyurl.com/socialeyes-beatlab. In the first part of this video example (tableau: “canjo”), the two performers are seated very near to each other. Even at this close distance, we can clearly distinguish audience-gaze on one or the other performer, vs. their instruments, and follow shifts in audience attention between these elements. When this “canjo” tableau ends, the two performers spread out and we see the gaze follow them around the stage. Once one of the performers starts speaking (tableau: “pod rattle incident”), the audience mostly maintain their focus on him, rarely glancing toward the other performer who is gently striking metallic objects. This video highlights that the transformed gaze reliably follows the two musicians around the stage and provides proof-of-concept that our homography approach successfully projects gaze from a large audience in challenging scenes.
@@
To study the gaze patterns over an evolving scene and compare gaze data between different viewers, we map gaze from each viewer’s frame of reference to the scene’s frame of reference with our Homography module (see Section 3.2.6). Figure 7 presents individual gaze heatmaps from all participants and the aggregate gaze activity from all participants for each session overlaid on a reference frame from that session. Projecting the gaze to the static centerview perspective gets rid of the independent positions and head movements of each participant, thereby getting rid of the central bias. Heatmaps of transformed gaze coordinates (Figure 7A) demonstrate higher gaze dispersion and multiple discrete clusters for each session duration. These clusters coincide with the presentation screen area for the film viewing sessions and the performers’ movement activity for the concert performance sessions (see averaged heatmaps for each session in Figure 7B). Higher gaze dispersion can be observed during performance viewing than during film viewing, on both days, which could be partially explained by the larger physical space covered on the stage by the performers during the concert, in comparison to the presentation screen size for the film.
Figure 8 highlights different trends in the position of the heatmap centroid (i.e., the peak heatmap intensity location; marked by the intersection of white dashed lines), based on seating position. Investigating the gaze heatmaps for different participant seating clusters, we observe that participants in the right half of the audience spent more time exploring the right side of the stage, reflected by larger (higher gaze dispersion) and more intense (higher gaze duration) clusters in the right half and vice versa (see Figure 8, panels A and C). This effect is, however, only observed in the performance sessions and not when the participants viewed the pre-recorded documentary film. The heatmaps could further be aggregated to observe similar effects with different participant clusters, such as clustering by participant columns (Figure 8A) or rows (Figure 8B).
@@ 4.2.3. Analysis
Further comparisons between individual heatmaps can be made with the help of metrics from our Analysis module (see Section 3.2.8). Pairwise SIM and CC scores were calculated for each participant pair for each of the four event sessions. For raw egocentric gaze, the mean SIM score during film viewing was 0.35 (SD = 0.21) for G1F and 0.38 (SD = 0.19) for G2F, while the mean SIM during the performance was 0.40 (SD = 0.18) for G1P and 0.43 (SD = 0.17) for G2P. The mean CC was 0.37 (SD = 0.30) for G1F, 0.42 (SD = 0.28) for G2P, 0.43 (SD = 0.26) for G1P and 0.46 (SD = 0.25) for G2P. When using homography-transformed gaze, the scores in all sessions, for both of these metrics, increase drastically (see Figure 9; left column), nearly doubling in most cases. A 2D chart of seat positions (row, column) for each participant (see Figures 6 and 7 for audience seating positions on each day) was used to calculate euclidean distances (DI) between each participant pair. Calculated SIM scores on the homography-transformed gaze were found to be negatively correlated with the euclidean distances during the performance session on both days (i.e., the closer two participants were to each other, the stronger was the correlation between their gaze patterns and vice versa). The Pearson correlation coefficient between SIM and DI was -0.10 (p = 0.039) for G1P and -0.17 (p ¡ 0.001) for G2P.
[Spatial gaze measures]Figure 9. Overall, the gaze entropy is higher and the gaze velocity is lower for performance sessions, as compared to the film sessions. However, further analysis is required to confirm these trends.
@@
The mean gaze entropy over all participants during film viewing was 0.76 (SD = 0.03) for G1F and 0.77 (SD = 0.03) for G2F. The mean entropy was higher when participants watched the performance, with a value of 0.82 (SD = 0.03) for G1P and 0.83 (SD = 0.03) for G2P. The higher values represent more spatial gaze dispersion when participants watch the performance as compared to the documentary film. The average gaze velocity (calculated in pixels/sample) during film viewing was found to be 22.99 (SD = 7.25) for G1P and 36.00 (SD = 5.90) for G2P. During the performance sessions, the average gaze velocity was 16.76 (SD = 6.79) for G1P and 22.11 (SD = 8.03) for G2P. Interestingly, gaze velocity was higher during film viewing on both days which signifies longer saccadic movements when viewing pre-recorded media on a screen, in comparison to a live concert performance.
[Blinks]Figure 10. The plots demonstrate individual as well as combined blink activity of the audience members over time, which can be directly compared with the audiovisual stimulus presented at the moment.
@@ 4.2.4. Exploratory Temporal Analysis
[Temporal metrics]Figure 11. Fully described in the text.
[Gaze behaviour during concert viewing]Figure 12. Fully described in the text.
The synchronised gaze data from multiple devices mapped to a common coordinate space (centralview) further facilitates novel spatiotemporal investigations of collective gaze measures over time. In Figure 11, we visualise the measures of standard deviation, minimum contour area, and points in frame, over the entire duration of a film-viewing session (1 hr 20 min film, plus some minutes before/after) from 30 people (G1). These measures–computed by the Analysis module (see Section 3.2.8)–reveal collective gaze dynamics over time. Standard deviation of the gaze distribution of all participants at any given time is represented in the horizontal and vertical directions by the ”SD x” and ”SD y” measures (top and second panels, respectively). The ”Normalised contour area” (third panel) represents overall gaze dispersion as the area of a minimum convex hull enclosing gaze from all participants. The number of gaze points falling within the centralview scene at a given timepoint is represented by the ”Points in Frame” measure (bottom panel). The time series of all four measures highlight clear event boundaries such as the start and stop of the film presentation (vertical dotted lines). Future analyses will explore which features of the film drive changes in these collective gaze dynamics.
In another example, visualising the Normalised contour area measure for the entire concert performance (Figure 12) shows similar collective gaze behaviour from audiences during each performance tableau (represented by vertical colored lines) for each participant group. These collective gaze measures and corresponding time series could be further analysed to investigate temporal interactions with the audiovisual stimulus features (outside the scope of the current analysis). We provide these examples here to highlight the rich and vast amount of possibilities from shared gaze metrics in dynamic social settings.
@@ 5. General Discussion
Results from our utility test demonstrate the end-to-end implementation of the recording mode of SocialEyes in a real-world application. SocialEyes tackles 1) critical challenges of time synchronisation of multi-sensor data to scale data collection, 2) monitoring and troubleshooting to ensure reliable data collection over long durations, 3) semantic gaze mapping of the synchronised gaze data to analyse collective gaze patterns, and finally, 4) shared gaze visualisation and analysis to achieve novel insights from the collected data. The proposed framework aims to reduce technological and logistical barriers in conducting naturalistic eye-tracking experiments with large social groups, allowing replication/refutation and extension of previous lab studies in social, naturalistic contexts (Kulke and Hinrichs, 2021; Laidlaw et al., 2016; Foulsham and Kingstone, 2017).
@@ 5.1.1. Edge-cases in gaze projection
Future objectives for our framework would be to further investigate and optimise gaze mapping accuracy in different environments. In our previous pilot study (Saxena and Fink, 2023), we identified a few edge case scenarios that pose a challenge to the homography transformations. In summary, the SuperGlue + RANSAC approach demonstrated robustness to head movements, illumination changes, and scene occlusions, as observed in our current utility study as well. However, close distances between the observer and scene as well as duplicate objects in a scene can result in false predictions. Systematically identifying these edge failure cases in challenging environments and addressing them—either through software and algorithmic updates or by establishing well-defined best practices for experimental setups—could significantly improve the robustness and reliability of gaze projections. To automate the collection of ground-truth gaze locations in these investigations, fiducial markers can be used by guiding participants to fixate on specific targets–as demonstrated in Appendix A.
@@ 5.1.2. Alternatives to planar homography
The assumption of a substantial distance between the observer(s) and the scene is a theoretical limitation of the current homography-based gaze projection as it projects a plane at sufficiently large (infinite) distance from one view to another. This assumption allows for more reliable and calibration-free projections, particularly when dealing with moving cameras, such as the head-mounted egoview, and dynamic (non-static) scenes, such as a music performance. We show the validity of this assumption for our current setup in Appendix A. Future optimisations can also consider multi-view homography where the scene is captured by multiple centralview cameras and each egoview camera is related with the centralview corresponding to its position. This strategy would enable more precise cross-view analyses and spatial consistency in wide-view environments. To eliminate the 2D plane assumption, two alternative approaches are possible. Given the intrinsic parameters of the two cameras (egoview and centralview) and a common 3D coordinate system, a 2D point from the egoview can be projected to the 3D world coordinate system and then to the 2D centralview. This approach relies on pre-existing information about the locations of objects in the 3D world space, which is highly unfeasible, particularly in dynamic scenes. The second approach involves calibrating extrinsic positions of each egoview-centralview camera pair and projecting the 2D point from one camera’s view to another using epipolar geometry. This triangulation approach provides a general (ideal) solution, particularly for scenes with varying depths or close viewing distances where the homography planar assumption is not valid. However, it relies on accurate calibration parameters between the two cameras which is challenging with non-stationary cameras such as the head-mounted egoview cameras. As a result, the application of this approach relies on accurate and frequent re-calibration of the egoview cameras and prominent (unobstructed, well-lighted, strategically placed) visual markers that can be detected in the dynamic scene. These requirements might often interfere with the experimental setup, since the control of lighting and marker placement is not always possible. Additionally, markers create a visual distraction for viewers that would interfere with scientific investigations of visual attention. As an alternative to visual markers, localisation of the egoview camera (and therefore the eye-tracker) is also possible with invisible infrared markers (Dogan et al., 2022), feature-based tracking via algorithms such as visualSLAM (Chen et al., 2022) and optical flow (Beauchemin and Barron, 1995), marker-less methods (Jin et al., 2024), or using the on-device 9DOF Inertial Measurement Unit (IMU) sensors. However, each method has limitations, such as marker detection robustness over distance, the need for a pre-registered 3D world model in model-based approaches, and the reliability of the IMU sensor with the existing eye-tracker (see Section 5). Overall, triangulation based methods might offer benefits, such as depth compensation and runtime efficiency, warranting future investigations to evaluate their trade-offs in performance, accuracy, and ecological validity compared to the current homography approach.